# On prior smoothing with discrete spatial data in the context of disease mapping

### v1

@@ 2.1.1 Model Fitting
To complete model specification, the following hyperpriors were used. Usual choices for the intercept $\alpha$ and standard deviations $\sigma$ are flat priors and vague uniform priors, respectively. These choices are maintained for all priors defined above. The specific hyperparameters defined for each prior are $\eta\sim U(-1,1)$ (following instructions of NIMBLE user manual (de Valpine et al., 2024)), $\lambda\sim U(0,1)$ and for the range parameter $\psi$ of the exponential correlation function, we employ uniform priors with support allowing ranges up to the maximum interpoint distance over the region (Wang and Gelfand, 2014).
@@ 3 The Poisson-Gamma model
So, we can obtain explicitly the value of any of our proposed empirical metrics. We can immediately see that, as $\sigma_{r}^{2}\rightarrow\infty$, no smoothing results, all of our metrics are equal to 00. Consequently, what is of interest for a given map, is the behavior of the smoothing as $\sigma_{r}^{2}\rightarrow\infty$ (given $\mu_{r}$). Here, at $\sigma_{r}^{2}=0$, Equation 2 becomes $\mu_{r}-\hat{r}_{i}$, providing the maximum smoothing. With a set of $\{O_{i}\}$, we can quantify the empirical smoothing associated with the map.
@@ 4.1 Within prior simulation study
Here, we consider the seven priors proposed in subsection 2.1 individually: an independent (non)spatial prior, the five different CAR priors, and the GP prior using a covariance function employing the distance between centroids. More precisely, we compare theoretical smoothing with empirical smoothing to assess whether the model’s smoothing, as evaluated using the TCV metric aligns with that arising through the empirical metrics. To accomplish our objective, we employed Scenario 2, as described in subsection 4.2 focusing on the spatial structure of peninsular Spain with varying levels of disaggregation. Specifically, we present results for a disaggregation level of $A=47$, corresponding to the provinces of continental Spain, and then increase the number of areas to $A=100$ and $A=300$. It is worth noting that similar studies were conducted in other scenarios, yielding consistent conclusions (but omitted here). It is also worth noting, and is intuitive, that increased disaggregation results in less smooth maps, hence increased prior smoothing.
@@
Table 2: Empirical smoothing criteria, MSS, RMSS,the maximum values, $maxMSS$ and $maxRMSS$, and smoothing proportion together with the theoretical smoothing metric, TCV, for the different parameter values of the iCAR priors.
@@
We start with the iCAR prior. The amount of smoothing induced by this prior is influenced only by the value of $\sigma^{2}$, with smoothing decreasing as $\sigma^{2}$ increases (refer to subsection 2.1). Five different options are considered for $\sigma^{2}$, ranging from $10^{-4}$ to $0.25$. Table 2 presents the empirical smoothing criteria values considered in this work (MSS, RMSS, $maxMSS$, $maxRMSS$ and the SP), alongside the theoretical smoothing criteria (TCV). Results indicate that theoretical smoothing aligns with empirical metrics, with smoothing decreasing as $\sigma^{2}$ increases. Additionally, empirical smoothing remains nearly constant for very small values of $\sigma^{2}$, while the theoretical smoothing criterion approaches zero. This suggests we have reached maximum smoothing, similar to the baseline case of the Poisson-Gamma model (see section 3). Conversely, as $\sigma^{2}$ values increase, smoothing continues to decrease, but the pace of decrease slows. In contrast, the theoretical measure consistently increases. Across different numbers of areas, higher empirical values (MSS, RMSS, and maximum values) observed with $A=300$ result in lower smoothing and lower SP values compared to $A=47$, complicating comparisons among spatial structures. Additionally, for a TCV value of zero, the SP is lower for $A=300$ than for $A=47$ indicating that the maximum smoothing decreases as the number of areas increases. Moreover, a faster increase in TCV is observed with a greater number of areas, suggesting that smoothing approaches an asymptote more quickly as the spatial disaggregation increases.
Table 3: Empirical smoothing criteria, MSS, RMSS, the maximum values, $maxMSS$ and $maxRMSS$, and the smoothing proportion (SP), together with the theoretical smoothing metric (TCV) for the different parameter values of the LCAR priors.
@@
For the LCAR prior, the level of smoothing induced is influenced by the values of $\sigma^{2}$ and the spatial dependence parameter $\lambda$. Therefore, in this simulation study we vary the $\sigma^{2}$ and $\lambda$ values. Specifically, we considered three values for $\lambda$: $0.1$, $0.5$, and $0.9$ with a full range of $\sigma^{2}$ values, from very small to very large. We present results for $\sigma^{2}=10^{-4},0.0025,0.0081,0.04,0.25$, to assess whether theoretical smoothing aligns with that determined through empirical metrics. Recall that, theoretically, as $\sigma^{2}$ increases, the expected smoothing decreases, whereas as $\lambda$ increases, the smoothing increases. Table 3 shows the empirical smoothing criteria, MSS, RMSS, the maximum values, $maxMSS$ and $maxRMSS$, and the smoothing proportion, together with the TCV for the different parameter values. The results show a strong alignment between theoretical and empirical metrics across both levels of spatial disaggregation. For $\sigma^{2}=10^{-4}$, empirical smoothing remains nearly constant regardless of changes in the $\lambda$ parameter, while the theoretical smoothing metric approaches zero, indicating that maximum smoothing has been achieved. In contrast, as $\sigma^{2}$ increases, the $\lambda$ parameter still affects the amount of smoothing, but the effect is almost negligible, whereas the TCV increases. The smoothing tends to an asymptote. Similar to the iCAR prior, the results indicate that the maximum smoothing decreases as the number of areas increases, and that smoothing reaches its asymptote more quickly as spatial disaggregation increases. Furthermore, higher TCV values are associated with lower SP values as the number of areas increases.
@@ 4.2 Across priors simulation study
To smooth the surfaces, we adopt the spatial covariance matrix, $\mathbf{C}(\boldsymbol{\theta})=\sigma_{C}^{2}\mathbf{R}(\boldsymbol{\varphi})$, with $\mathbf{R}(\boldsymbol{\varphi})$ a correlation matrix whose size corresponds to the number of grid points. The $(i,j)$ elements of $\mathbf{R}(\boldsymbol{\varphi})$ are computed using a Matérn correlation function,
@@
To compare the extent of smoothing across priors, we fit the Poisson-logitNormal model with independent (non)spatial priors, referred to as iid, along with GP prior and the five CAR priors iCAR, BYM, pCAR, LCAR, and BYM2 as defined in subsection 2.1, to each scenario. To compare the smoothness of the priors, we compute empirical smoothing metrics — the MSS, RMSS, maximum MSS, maximum RMSS and SP (as detailed in subsection 2.2) — for each prior model and scenario.
Table 4: MSS, RMSS, maximum MSS, maximum RMSS and SP values reached by each spatial prior model in Scenario 1 and Scenario 3, across the spatial regions of Spain and England. For Spain, results are presented for spatial disaggregation levels of $A=47$ and $A=100$.
@@
Table 4 presents the empirical smoothing metrics for each prior model, scenario and spatial structure. The results include the spatial structures of 47 provinces and 100 areas in Spain, as well as the spatial structure of England. Metrics for additional spatial structures are provided in Appendix C. Notable, for all spatial structures, Scenario 3 exhibits greater empirical smoothing criteria values. This scenario has greater spatial variability, compared with Scenario 1 which has a clear spatial pattern. Moreover, in comparing the smoothing metric values between spatial structures, an increase in smoothness is observed as the spatial region is divided into a greater number of areas. Specifically, for the two spatial structures of Spain, the smoothness increases approximately fivefold when the number of areas rises from 47 to 100, with a similar trend observed in both scenarios. This pattern continues as the number of areas increases to 300, with the smoothing increasing threefold from 100 to 300 areas (refer to Appendix C). Additionally, when comparing the spatial structures of Spain and England, both with approximately 100 areas, England exhibits higher smoothing metric values, indicating a greater level of smoothness.
@@
When analyzing the spatial structure in England, the CAR spatial priors exhibit similar empirical criteria values. For mean values (MSS and RMSS), these priors yield the lowest values, with the LCAR prior showing a slight improvement in smoothing. Conversely, the iid prior achieves the highest mean values. Regarding the maximum values, the iid prior records the lowest values, while the GP prior achieves the highest. In terms of proportions, all priors show similar values in Scenario 1, whereas Scenario 2 reveals disparities, with the LCAR prior exhibiting less smoothing. It is worth noting that consistent conclusions arise across all of the analyzed scenarios (some of which are omitted here for brevity). Additionally, the results show that decreasing the variability of the rates leads to higher smoothing criteria values.
@@ 5 Data illustration
We now turn to the analysis of real datasets. Specifically, we investigate the smoothing effects induced by the seven spatial priors proposed in subsection 2.1 to clarify the theoretical and simulation findings for these datasets. As observed in subsection 4.1, the $\sigma^{2}$ parameter significantly affects the amount of smoothing induced by the spatial priors and is a common parameter across all priors. Therefore, we first assume different informative prior distributions for the $\sigma^{2}$ parameter of the proposed spatial priors. We define three informative prior distributions: small-size $Unif(0,0.01)$, medium-size $Unif(0.01,0.16)$, and large-size $Unif(0.16,100)$ prior distributions. Additionally, we include a fourth prior distribution for the $\sigma^{2}$ parameter, a uniform prior distribution on $(0,1000)$.
We use two datasets representing two different spatial regions, as noted in the previous section. The first spatial region corresponds to peninsular Spain, divided into $A=47$ (the real provinces), $A=100$ and $A=300$. The second spatial region relates to 106 clinical commissioning groups of England. Both datasets include counts for lung cancer and the corresponding population at risk, focusing on aggregate data for females in Spain from 2019 to 2021 and for males in England in 2017. The Spanish Statistical Office (INE) provided the dataset for Spain. The cancer mortality data for England has been supplied by the National Cancer Registration and Analysis Service (NCRAS) and population data by the Office for National Statistics (ONS). Results for peninsular Spain with $A=47$ and $A=300$ are presented in this section, while results for Spain with $A=100$, along with the England spatial structure, are detailed in Appendix D.
@@ 5.1 Lung cancer deaths in Spain
When analyzing the posterior mean rates per 100,000 inhabitants obtained for each spatial prior, see Figure A7 and Figure A8 for $A=47$ and Figure A9 and Figure A10 for $A=300$ in Appendix D, a significant degree of smoothing is evident when using a small-size prior distribution across all spatial priors. Higher rates are concentrated in north-western Spain, gradually decreasing as we move outward from this hotspot. However, the lower rates initially observed in the central-southern provinces in the crude rates disappear, resulting in a smoother surface overall. This smoothing effect becomes more pronounced as the number of areas increase. For $A=47$, a similar smoothing effect is observed across all spatial priors for each variance prior distribution. When the number of areas in peninsular Spain is increased to $A=300$, the surfaces obtained with the different variance prior distributions become noticeably smoother compared to the crude rates, though differences among the spatial priors begin to emerge. With a small-size informative prior distribution, greater smoothing is observed, particularly with the iid prior, which produces similar rate estimates for most areas. The other spatial priors exhibit comparable rate distributions, with higher rates concentrated in Asturias and Cantabria that gradually decrease outward from this hotspot. Notably, with the GP prior, some disparities emerge, as higher rates are maintained in areas such as Zaragoza and along the Mediterranean coast. When assuming medium-size informative prior distributions, the disparities among the spatial priors diminish. However, for large-size informative prior distributions, the BYM prior produces a slightly less smooth surface. As mentioned earlier, the uniform prior presents very similar smoothing to that seen with the medium-size informative prior distribution for both spatial disaggregations.
In conclusion, it is worthwhile to illuminate the smoothing effects introduced by spatial priors when interpreting and drawing conclusions about the spatial distribution of a disease. This consideration becomes more useful when examining results using small-size informative prior distributions or when analyzing a larger number of areas.
@@ 6 Discussion
Little discussion has appeared in the literature with regard to quantification and comparison of how and how much various neighbor-based priors introduce spatial smoothing in the context of disease mapping modeling. This becomes of interest because model performance is not a question of goodness-of-fit. We have taken up that issue here through both simulation and real data. We have offered a theoretical metric for assessing the expected smoothing associated with a particular prior in terms of its parameters and we have offered empirical metrics to quantify the resulting smoothing achieved by a particular prior. The path here is not intended to assert that one prior is “better” than another. Rather, the goal is to enable comparison of smoothing within and across these priors.
@@
The utility of the analysis developed here is to better enable the practitioner working with areal disease data to appreciate the smoothing consequences of choice of disease mapping prior and specification of hyperpriors on the parameters of the prior. Maps can provide visual comparison of resultant smoothing but explicit quantification may offer a useful supplement. Further, since our methodology is generally applicable, such smoothing assessment under fitting of a particular disease mapping model specification could be useful in interpreting the results/implications of a disease mapping analysis.
@@
Future work could explore the dynamic disease mapping setting where some autoregressive specification is introduced into the modeling (Knorr-Held, 2000). Then, we might ask about the effect of temporal dependence on smoothing over time. Also, there is now a rich literature on multivariate disease mapping to capture interaction/association in disease risk and rates (Martínez-Beneito and Botella-Rocamora, 2019, Chapter 8). When within unit dependence is introduced into the joint modeling, how does such modeling affect smoothing compared with marginal modeling of the diseases?

### v2

@@ 2.1.1 Model Fitting
To complete model specification, the following hyperpriors were used. Usual choices for the intercept $\alpha$ and standard deviations $\sigma$ are flat priors and vague uniform priors, respectively. These choices are maintained for all priors defined above. The specific hyperparameter distributions defined for each prior are as follows: the spatial dependence parameter $\eta\sim U(-1,1)$ (following instructions of NIMBLE user manual (de Valpine et al., 2024)); the spatial dependence parameter $\lambda\sim U(0,1)$; and for the range parameter $\psi$ of the exponential correlation function, we employ uniform priors with support allowing ranges up to the maximum interpoint distance over the region (Wang and Gelfand, 2014).
@@ 3 The Poisson-Gamma model
So, we can obtain explicitly the value of any of our proposed empirical metrics. We can immediately see that, as $\sigma_{r}^{2}\rightarrow\infty$, no smoothing results, all of our metrics are equal to 0. Consequently, what is of interest for a given map, is the behavior of the smoothing as $\sigma_{r}^{2}\rightarrow\infty$ (given $\mu_{r}$). Here, at $\sigma_{r}^{2}=0$, Equation 2 becomes $\mu_{r}-\hat{r}_{i}$, providing the maximum smoothing. With a set of $\{O_{i}\}$, we can quantify the empirical smoothing associated with the map.
@@ 4.1 Within prior simulation study
Here, we consider the seven priors proposed in subsection 2.1 individually: an independent (non)spatial prior, the five different CAR priors, and the GP prior using a covariance function employing the distance between centroids. More precisely, we compare theoretical smoothing with empirical smoothing to assess whether the model’s smoothing, as evaluated using the TCV metric aligns with that arising through the empirical metrics. To accomplish our objective, we employed Scenario 2, as described in subsection 4.2 focusing on the spatial structure of peninsular Spain with varying levels of disaggregation. Specifically, we present results for a disaggregation level of $A=47$, corresponding to the provinces of peninsular Spain, and then increase the number of areas to $A=100$ and $A=300$. It is worth noting that similar studies were conducted in other scenarios, yielding consistent conclusions (but omitted here). It is also worth noting, and is intuitive, that increased disaggregation results in less smooth maps, hence increased prior smoothing.
@@
We start with the iCAR prior. The amount of smoothing induced by this prior is influenced only by the value of $\sigma^{2}$, with smoothing decreasing as $\sigma^{2}$ increases (refer to subsection 2.1). Five different options are considered for $\sigma^{2}$, ranging from $10^{-4}$ to 0.250.250.25. Table 2 presents the empirical smoothing criteria values considered in this work (average of MSS and RMSS, $maxMSS$, $maxRMSS$ and the SP), alongside the theoretical smoothing criteria (TCV). Results indicate that theoretical smoothing aligns with empirical metrics, with smoothing decreasing as $\sigma^{2}$ increases. Additionally, empirical smoothing remains nearly constant for very small values of $\sigma^{2}$, while the theoretical smoothing criterion approaches zero. This suggests we have reached maximum smoothing, similar to the baseline case of the Poisson-Gamma model (see section 3). Conversely, as $\sigma^{2}$ values increase, smoothing continues to decrease, but the pace of decrease slows. In contrast, the theoretical measure consistently increases. Across different numbers of areas, higher empirical values (MSS, RMSS, and maximum values) observed with $A=300$ result in lower smoothing and lower SP values compared to $A=47$, complicating comparisons among spatial structures. Additionally, for a TCV value of zero, the SP is lower for $A=300$ than for $A=47$ indicating that the maximum smoothing decreases as the number of areas increases. Moreover, a faster increase in TCV is observed with a greater number of areas, suggesting that smoothing approaches an asymptote more quickly as the spatial disaggregation increases.
Table 2: Empirical smoothing criteria, the average values of MSS, RMSS,the maximum values, $maxMSS$ and $maxRMSS$, and smoothing proportion together with the theoretical smoothing metric, TCV, for the different parameter values of the iCAR priors.
@@
Table 3: Empirical smoothing criteria, the average values of MSS, RMSS, the maximum values, $maxMSS$ and $maxRMSS$, and the smoothing proportion (SP), together with the theoretical smoothing metric (TCV) for the different parameter values of the LCAR priors.
@@
For the LCAR prior, the level of smoothing induced is influenced by the values of $\sigma^{2}$ and the spatial dependence parameter $\lambda$. Therefore, in this simulation study we vary the $\sigma^{2}$ and $\lambda$ values. Specifically, we considered three values for $\lambda$: 0.10.10.1, 0.50.50.5, and 0.90.90.9 with a full range of $\sigma^{2}$ values, from very small to very large. We present results for $\sigma^{2}=10^{-4},0.0025,0.0081,0.04,0.25$, to assess whether theoretical smoothing aligns with that determined through empirical metrics. Recall that, theoretically, as $\sigma^{2}$ increases, the expected smoothing decreases, whereas as $\lambda$ increases, the smoothing increases. Table 3 shows the empirical smoothing criteria, the average values of MSS and RMSS, the maximum values, $maxMSS$ and $maxRMSS$, and the smoothing proportion, together with the TCV for the different parameter values. The results show a strong alignment between theoretical and empirical metrics across both levels of spatial disaggregation. For $\sigma^{2}=10^{-4}$, empirical smoothing remains nearly constant regardless of changes in the $\lambda$ parameter, while the theoretical smoothing metric approaches zero, indicating that maximum smoothing has been achieved. In contrast, as $\sigma^{2}$ increases, the $\lambda$ parameter still affects the amount of smoothing, but the effect is almost negligible, whereas the TCV increases. The smoothing tends to an asymptote. Similar to the iCAR prior, the results indicate that the maximum smoothing decreases as the number of areas increases, and that smoothing reaches its asymptote more quickly as spatial disaggregation increases. Furthermore, higher TCV values are associated with lower SP values as the number of areas increases.
@@ 4.2 Across priors simulation study
[…]
@@
To compare the extent of smoothing across priors, we fit the Poisson-logitNormal model with independent (non)spatial priors, referred to as iid, along with GP prior and the five CAR priors iCAR, BYM, pCAR, LCAR, and BYM2 as defined in subsection 2.1, to each scenario. To compare the smoothness of the priors, we compute empirical smoothing metrics — the average values of MSS and RMSS, maximum MSS, maximum RMSS and SP (as detailed in subsection 2.2) — for each prior model and scenario.
Table 4: Average values of MSS, RMSS, maximum MSS, maximum RMSS and SP values reached by each spatial prior model in Scenario 1 and Scenario 3, across the spatial regions of Spain and England. For Spain, results are presented for spatial disaggregation levels of $A=47$ and $A=100$.
@@
Table 4 presents the empirical smoothing metrics for each prior model, scenario and spatial structure. The results include the spatial structures of 47 provinces and 100 areas in Spain, as well as the spatial structure of England. Metrics for additional spatial structures are provided in Appendix C. Notable, for all spatial structures, Scenario 3 exhibits greater empirical smoothing criteria values. This scenario has greater spatial variability, compared with Scenario 1 which has a clear spatial pattern. Moreover, in comparing the smoothing metric values between spatial structures, an increase in smoothness is observed as the spatial region is divided into a greater number of areas. Specifically, for the two spatial structures of Spain, the smoothness increases approximately fivefold when the number of areas rises from 47 to 100, with a similar trend observed in both scenarios. This pattern continues as the number of areas increases to 300, with the smoothing increasing threefold from 100 to 300 areas (refer to Appendix C). Additionally, when comparing the spatial structures of Spain and England, both with approximately 100 areas, England generally exhibits higher MSS and RMSS values, but lower maximum smoothing values and a smaller proportion of smoothing.
@@
When analyzing the spatial structure in England, the CAR spatial priors exhibit similar empirical criteria values. For mean values (MSS and RMSS), these priors yield the lowest values, with the LCAR prior showing a slight improvement in smoothing. Conversely, the iid prior achieves the highest mean values. Regarding the maximum values, the iid prior records the lowest values, while the GP prior achieves the highest. In terms of proportions, all priors show similar values in Scenario 1, whereas Scenario 3 reveals disparities, with the LCAR prior exhibiting less smoothing. It is worth noting that consistent conclusions arise across all of the analyzed scenarios (some of which are omitted here for brevity). Additionally, the results show that decreasing the variability of the rates leads to higher smoothing criteria values.
@@ 5 Data illustration
We now turn to the analysis of real datasets. Specifically, we investigate the smoothing effects induced by the seven spatial priors proposed in subsection 2.1 to clarify the theoretical and simulation findings for these datasets. As observed in subsection 4.1, the $\sigma^{2}$ parameter significantly affects the amount of smoothing induced by the spatial priors and is a common parameter across all priors. Therefore, we first assume different informative prior distributions for the $\sigma^{2}$ parameter of the proposed spatial priors. We define three informative prior distributions: small-size $Unif(0,0.01)$, medium-size $Unif(0.01,0.16)$, and large-size $Unif(0.16,100)$ prior distributions. Additionally, we include a fourth prior distribution for the $\sigma^{2}$ parameter, a uniform prior distribution on (0,1000)(0,1000)( 0 , 1000 ).
We use two datasets representing two different spatial regions, as noted in the previous section. The first spatial region corresponds to peninsular Spain, divided into $A=47$ (the real provinces), $A=100$ and $A=300$. The second spatial region relates to 106 clinical commissioning groups of England. Both datasets include counts for lung cancer and the corresponding population at risk, focusing on aggregate data for females in Spain from 2019 to 2021 and for males in England in 2017. The Spanish Statistical Office (INE) provided the dataset for Spain. The cancer mortality data for England has been supplied by the National Cancer Registration and Analysis Service (NCRAS) and population data by the Office for National Statistics (ONS). Results for peninsular Spain with $A=47$ and $A=300$ are presented in this section, while results for Spain with $A=100$, along with the England spatial structure, are detailed in Appendix C.
@@ 5.1 Lung cancer deaths in Spain
When analyzing the posterior mean rates per 100,000 inhabitants obtained for each spatial prior, see Figure 10 and Figure 11 for $A=47$ and Figure 12 and Figure 13 for $A=300$ in Appendix D, a significant degree of smoothing is evident when using a small-size prior distribution across all spatial priors. Higher rates are concentrated in north-western Spain, gradually decreasing as we move outward from this hotspot. However, the lower rates initially observed in the central-southern provinces in the crude rates disappear, resulting in a smoother surface overall. This smoothing effect becomes more pronounced as the number of areas increase. For $A=47$, a similar smoothing effect is observed across all spatial priors for each variance prior distribution. When the number of areas in peninsular Spain is increased to $A=300$, the surfaces obtained with the different variance prior distributions become noticeably smoother compared to the crude rates, though differences among the spatial priors begin to emerge. With a small-size informative prior distribution, greater smoothing is observed, particularly with the iid prior, which produces similar rate estimates for most areas. The other spatial priors exhibit comparable rate distributions, with higher rates concentrated in Asturias and Cantabria that gradually decrease outward from this hotspot. Notably, with the GP prior, some disparities emerge, as higher rates are maintained in areas such as Zaragoza and along the Mediterranean coast. When assuming medium-size informative prior distributions, the disparities among the spatial priors diminish. However, for large-size informative prior distributions, the BYM prior produces a slightly less smooth surface. As mentioned earlier, the uniform prior presents very similar smoothing to that seen with the medium-size informative prior distribution for both spatial disaggregations.
In conclusion, the behavior observed in the simulation studies is consistent when applied to real data sets, confirming that the choice of spatial prior influences the level of smoothing in the estimates. By analyzing the Spain and England datasets, each with distinct spatial structures, we also assess the robustness of our findings across different geographical contexts. Two main conclusions can be reached from the real data analyses. First, as the number of spatial units increases, the effect of smoothing becomes more pronounced, underscoring the need for caution when interpreting disease rates in large-area settings, where oversmoothing may obscure meaningful spatial variation. Second, the choice of prior distribution for the hyperparameters of the spatial model influences the level of smoothing, with different specifications potentially masking important spatial patterns. These findings highlight the practical importance of understanding the smoothing effects introduced by spatial priors, which is essential for accurately interpreting and drawing conclusions about the spatial distribution of a disease.
@@ 6 Discussion
Little discussion has appeared in the literature with regard to quantification and comparison of how and how much various neighbor-based priors introduce spatial smoothing in the context of disease mapping modeling. This becomes of interest because model performance is not a question of goodness-of-fit. In this work, we address this gap through both simulation studies and real data applications, proposing metrics to evaluate spatial smoothing from both theoretical and empirical perspectives. We introduce the Total Conditional Variance (TCV), defined as the sum of the conditional variances across all areal units, to evaluate the expected level of smoothing associated with a given spatial prior, i.e., this metric allows for a comparison of the smoothness induced by different hyperprior distributions within that spatial prior. It also provides insight into the theoretical behavior of smoothing in the Poisson-logitNormal model. Notably, we observe that the smoothing effect reaches a maximum and then, when $\sigma^{2}$ increases, it approaches zero—behavior that mirrors the properties of the Poisson-Gamma model. On the empirical side, we propose additional metrics to quantify the actual smoothing achieved by different priors in practice. These empirical measures allow us to confirm that model behavior aligns with theoretical expectations. Furthermore, they facilitate meaningful comparisons across different spatial priors in terms of their smoothing effects.
The within prior simulation study demonstrates that theoretical expectations are consistent with empirical performance, suggesting that $\sigma^{2}$ consistently exercises control over the expected level of smoothing across all spatial priors, even when additional parameters are present. The across prior simulation study further reveals that the choice of spatial prior plays an important role in the amount of smoothing achieved within a data-set. This effect becomes more pronounced as the number of areas increases. Furthermore, higher levels of smoothing are often observed when increasing the number of areas. These conclusions are further supported by the findings from the real data analyses. These analyses also demonstrate that, beyond the spatial prior defined, the selection of hyperprior distributions is also important, leading to varying degrees of smoothing within the priors.
When comparing the level of smoothing across spatial priors, it is not possible to draw a general conclusion, as the results vary depending on the spatial structure of the regions. This variability is evident in both the simulation studies and the real data analyses. However, across both spatial regions used in this work, the GP prior generally produces the highest levels of smoothing. In the case of England, a slight reduction in smoothing is observed when using the LCAR prior, consistent in both the real data and the simulations. For Spain, differences among the spatial priors become more pronounced as the number of areas increases, with the BYM prior exhibiting lower levels of smoothing in such settings.
@@
In conclusion, the utility of the analysis developed here is to better enable the practitioner working with areal disease data to appreciate the smoothing consequences of choice of disease mapping prior and specification of hyperpriors on the parameters of the prior. Maps can provide visual comparison of resultant smoothing but explicit quantification may offer a useful supplement. Further, since our methodology is generally applicable, such smoothing assessment under fitting of a particular disease mapping model specification could be useful in interpreting the results/implications of a disease mapping analysis.
@@
Future work could explore the dynamic disease mapping setting where some autoregressive specification is introduced into the modeling (Knorr-Held, 2000). Then, we might ask about the effect of temporal dependence on smoothing over time. Also, there is now a rich literature on multivariate disease mapping to capture interaction/association in disease risk and rates (Mart´ınez-Beneito and Botella-Rocamora, 2019, Chapter 8). When within unit dependence is introduced into the joint modeling, how can we formally assess resultant smoothing and how does such modeling affect smoothing compared with marginal modeling of the diseases?